# Continuous-time reinforcement learning for optimal switching over multiple regimes

### v1

@@ 2 Classical Optimal Switching Problem
$dX^{t,x,i,\alpha}_{s}=\mu(s,X^{t,x,i,\alpha}_{s},I_{s}^{t,i})ds+\sigma(s,X^{t,x,i,\alpha}_{s},I_{s}^{t,i})dW_{s},\quad s\in(t,T].$ (2.2)
@@ 3 Exploratory Formulation under Entropy Regularization
Given $M>0$, consider a smooth and non-decreasing cut-off function $\phi_{M}$ such that $\phi_{M}(x)=e^{x}$ for $x\leq M$, $\phi_{M}(x)\leq e^{x}$ for $x\in(M,M+1)$ and $\phi_{M}(x)=e^{M+1}$ for $x\geq M+1$. Hence, $\phi_{M}$ is bounded and Lipschitz continuous. Denote ${\cal D}_{N}:=\{(t,x):(t,x)\in{\cal D},|x|<N\}$. First, we will solve the following partial differential equation (PDE) systems: for $i\in\mathbb{I}_{m}$,
$\begin{cases}\displaystyle\frac{\partial V_{i}^{M,N}(t,x)}{\partial t}+\mathcal{L}^{i}_{x}V_{i}^{M,N}(t,x)+f(t,x,i)+\lambda\sum_{j\neq i}\phi_{M}\left(\frac{V^{M,N}_{j}(t,x)-V^{M,N}_{i}(t,x)-g_{ij}}{\lambda}\right)=0,\\[10.00002pt] \displaystyle\hfill(t,x)\in{\cal D}_{N},\\[10.00002pt] \displaystyle V_{i}^{M,N}(t,x)=K(T-t)+h(x),\quad(t,x)\in\partial{\cal D}_{N},\end{cases}$ (3.7)
@@
Consequently, we can extract from $\{V_{i}^{N}(t,x)\}$ a subsequence converging uniformly in $\overline{{\cal D}}$ together with the first $x$ , $t$-derivatives and second $x$-derivatives to a limit function $V_{i}^{\lambda}$, which is a solution to the HJB system (3.4). The uniqueness of the solution follows from Theorem 1.3 in kusano1965first. Thus, we complete the proof of the theorem. ∎
@@
$V_{i}^{\lambda}(t,x)\geq\mathbb{E}_{t,x,i}\bigg[\int^{s}_{t}f(\ell,X_{\ell},I_{\ell})ds-\sum_{k=1}^{\infty}g_{\kappa_{k-1}\kappa_{k}}{\bf 1}_{\{\tau_{k}\leq s\}}+\lambda\int^{s}_{t}R({\bm{\pi}}_{\ell},I_{\ell})ds+V_{I_{s}}^{\lambda}(s,X_{s})\bigg]$ (3.17)
@@ 4 Policy Iteration and Convergence
$\begin{cases}\displaystyle\frac{\partial V_{i}^{n}(t,x)}{\partial t}+\mathcal{L}^{i}_{x}V_{i}^{n}(t,x)+f(t,x,i)+H_{i}({\bm{\pi}}^{n}_{i}(t,x),V_{1}^{n}(t,x),\cdots,V_{m}^{n}(t,x))=0,\\ \displaystyle V_{i}^{n}(T,x)=h(x),\end{cases}$ (4.1)
@@
For $n\geq 1$, let $\Delta_{i}^{n}(t,x):=V_{i}^{n+1}(t,x)-V_{i}^{n}(t,x)$, for $i\in\mathbb{I}_{m}$ and $(t,x)\in\overline{{\cal D}}$. By using (4.1), $\Delta_{i}^{n}(t,x)$ satisfies
$\frac{\partial\Delta_{i}^{n}(t,x)}{\partial t}+\mathcal{L}^{i}_{x}\Delta_{i}^{\lambda}(t,x)+$ $H_{i}({\bm{\pi}}^{n+1}_{i}(t,x),V_{1}^{n+1}(t,x),\cdots,V_{m}^{n+1}(t,x))$
- $H_{i}({\bm{\pi}}^{n}_{i}(t,x),V_{1}^{n}(t,x),\cdots,V_{m}^{n}(t,x))=0,\quad\text{for}\penalty 10000\ (t,x)\in{\cal D},$ (4.4)
with the terminal condition $\Delta_{i}^{n}(T,x)=0$ for $x\in\mathds{R}^{n}$. From (4.3), we can see
${\bm{\pi}}_{i}^{n+1}(t,x)=\underset{{\bm{\pi}}_{i}}{\mathrm{arg\,max\,}}H_{i}({\bm{\pi}}_{i},V_{1}^{n}(t,x),\cdots,V_{m}^{n}(t,x)).$ (4.5)
It follows from (4) and (4.5) that, for $(t,x)\in{\cal D}$,
$\frac{\partial\Delta_{i}^{n}(t,x)}{\partial t}+\mathcal{L}^{i}_{x}\Delta_{i}^{n}(t,x)+\sum_{j\neq i}\pi_{ij}^{n+1}(t,x)\Delta_{j}^{n}(t,x)-\sum_{j\neq i}\pi_{ij}^{n+1}(t,x)\Delta_{i}^{n}(t,x)$
$=-H_{i}({\bm{\pi}}_{i}^{n+1}(t,x),V_{1}^{n+1}(t,x),\cdots,V_{m}^{n+1}(t,x))-\sum_{j\neq i}\pi_{ij}^{n+1}(t,x)(\Delta_{j}^{n}(t,x)-\Delta_{i}^{n}(t,x))$
$\quad+H_{i}({\bm{\pi}}^{n}_{i}(t,x),V_{1}^{n}(t,x),\cdots,V_{m}^{n}(t,x))$
$=H_{i}({\bm{\pi}}^{n}_{i}(t,x),V_{1}^{n}(t,x),\cdots,V_{m}^{n}(t,x))-H_{i}({\bm{\pi}}^{n+1}_{i}(t,x),V_{1}^{n}(t,x),\cdots,V_{m}^{n}(t,x))$
@@
By applying Theorem 1.3 in kusano1965first, we deduce that $\Delta^{n}_{i}(t,x)\geq 0$, that is, $V_{i}^{n+1}(t,x)\geq V_{i}^{n}(t,x)$, for all $i\in\mathbb{I}_{m}$ and $(t,x)\in\overline{{\cal D}}$.
On the other hand, for $n\geq 1$, let $\tilde{\Delta}_{i}^{n}(t,x):=V_{i}^{\lambda}(t,x)-V_{i}^{n}(t,x)$, for $i\in\mathbb{I}_{m}$ and $(t,x)\in\overline{{\cal D}}$. In a similar fashion, it can be shown that, for $(t,x)\in{\cal D}$,
$\frac{\partial\tilde{\Delta}_{i}^{n}(t,x)}{\partial t}+\mathcal{L}^{i}_{x}\tilde{\Delta}_{i}^{n}(t,x)+\sum_{j\neq i}\pi_{ij}^{n+1}(t,x)\tilde{\Delta}_{j}^{n}(t,x)-\sum_{j\neq i}\pi_{ij}^{n+1}(t,x)\tilde{\Delta}_{i}^{n}(t,x)$
$=H_{i}({\bm{\pi}}^{n}_{i}(t,x),V_{1}^{\lambda}(t,x),\cdots,V_{m}^{\lambda}(t,x))-H_{i}({\bm{\pi}}^{*}_{i}(t,x),V_{1}^{\lambda}(t,x),\cdots,V_{m}^{\lambda}(t,x))$
$\leq 0,$ (4.7)
and $\tilde{\Delta}_{i}^{n}(T,x)=0$ for $x\in\mathds{R}^{n}$. By applying Theorem 1.3 in kusano1965first again, $\tilde{\Delta}_{i}^{n}(t,x)\geq 0$, i.e., $V_{i}^{\lambda}(t,x)\geq V_{i}^{n}(t,x)$, for all $i\in\mathbb{I}_{m}$ and $(t,x)\in\overline{{\cal D}}$, which then completes the proof. ∎
@@
$\left|H_{i}({\bm{\pi}}_{i}^{n+1},V_{1}^{\lambda},\cdots,V_{m}^{\lambda})-H_{i}({\bm{\pi}}_{i}^{*},V_{1}^{\lambda},\cdots,V_{m}^{\lambda})\right|$
$\leq\left|\sum_{j\neq i}\pi_{ij}^{n}(V_{j}^{\lambda}-g_{ij}-V_{i}^{\lambda})+\lambda\sum_{j\neq i}(\pi_{ij}^{n}-\pi_{ij}^{n}\log\pi_{ij}^{n})\right|$
$\quad+\left|\left(\sum_{j\neq i}\pi_{ij}^{*}(V_{j}^{\lambda}-g_{ij}-V_{i}^{\lambda})+\lambda\sum_{j\neq i}(\pi_{ij}^{*}-\pi_{ij}^{*}\log\pi_{ij}^{*})\right)\right|$
@@
and $w_{i}^{n}(T,x)\geq 0$ for $x\in\mathds{R}^{n}$. By virtue of Theorem 1.3 in kusano1965first, we deduce $w^{n}_{i}(t,x)\geq 0$. That is,
@@
$F^{n}(t)\leq\frac{(C^{*})^{n}T^{n}}{n!}F^{1}(t),\quad\forall t\in[0,T].$ (4.14)
Because $F^{1}(t)$ is bounded, let $C_{1}=C^{*}T$ and $C_{2}=\sup_{t\in[0,T]}F^{1}(t)$. Then we obtain that desired result. ∎
@@
It follows from (3.13) and Assumption 2.2-(ii) that
$|V_{i}^{\lambda}(t,x)|\leq K_{f,h}(1+T)+\lambda\sup_{i\in\mathbb{I}_{m}}\left(\sum_{j\neq i}\exp\left(-\frac{g_{ij}}{\lambda}\right)\right)T\leq K_{f,h}(1+T)+\lambda(m-1)T$
@@
On the other hand, it follows from the definition of upper and lower weak limits that $\overline{V}_{i}(t,x)\geq\underline{V}_{i}(t,x)$, for any $i\in\mathbb{I}_{m}$ and $(t,x)\in\overline{{\cal D}}$. Thus, $\overline{V}_{i}(t,x)=\underline{V}_{i}(t,x)$, then denotes by
$V^{*}_{i}(t,x)=\overline{V}_{i}(t,x)=\underline{V}_{i}(t,x)\quad\text{for}\penalty 10000\ i\in\mathbb{I}_{m},(t,x)\in\overline{{\cal D}}.$
@@ 5 Reinforcement Learning Algorithm
Recall that given a feedback strategy ${\bm{\pi}}(t,x)=(\pi_{ij}(t,x))_{i,j\in\mathbb{I}_{m}}$, then the corresponding value function $(v_{1}^{{\bm{\pi}}},\cdots,v_{m}^{{\bm{\pi}}})$ satisfies the PDE system that for $i\in\mathbb{I}_{m}$,
@@
where the the Hamiltomian $H_{i}$ is given by (4.2). For simplicity, we omit the superscript $\bm{\pi}$}} and denote the value function as:
@@
and denote by $I=(I_{t}){t\geq 0}$ a continuous-time finite-state Markov chain with generator ${\bm{\pi}}=(\pi^{ij})_{i,j\in\mathbb{I}_{m}}$. Let us introduce the process $M=(M_{t})_{t\in[0,T]}$ given by
@@
Let ${\bm{\pi}}(t,x)=(\pi_{ij}(t,x))_{i,j\in\mathbb{I}_{m}}$ be a feedback strategy and $v(t,x,i)$ be the corresponding value function given by (5.2). Then the process $M=(M_{t})_{t\in[0,T]}$ given by (5.3) is a squar-intergrable martingale.
@@
$=M_{t^{\prime}}+\mathbb{E}\left[\int_{t^{\prime}}^{t}\int_{t^{\prime}}^{t}(D_{x}v(s,X_{s},I_{s}))^{\top}\sigma(s,X_{s},I_{s})dW_{s}\Big|\mathcal{F}_{t^{\prime}}\right]=M_{t^{\prime}}.$ (5.5)
@@
$\sup_{j\in\mathbb{I}_{m}}|v^{\xi_{i}}(\cdot,j)-v(\cdot,j)|_{C^{0}({\overline{{\cal D}}})}\leq\sup_{j\in\mathbb{I}_{m}}|v(\cdot,j)-v^{\xi^{*}}(\cdot,j)|_{C^{0}({\overline{{\cal D}}})}+\frac{C}{\epsilon^{\rho_{\xi}/2}}i^{-\frac{\nu\rho_{\xi}}{2}}.$ (5.14)
@@
$|\xi_{i}-\xi^{*}|^{2}\leq C\epsilon^{-\frac{1}{2}}i^{-\frac{\nu}{2}}$
@@
$\leq\sup_{j\in\mathbb{I}_{m}}|v^{\xi^{*}}(\cdot,)-v(\cdot,j)|_{C^{0}({\overline{{\cal D}}})}+\frac{C}{\epsilon^{\rho_{\xi}/2}}i^{-\frac{\nu\rho_{\xi}}{2}}.$
@@
Theorem 5.4 establishes a comprehensive error analysis for Algorithm 1, providing both theoretical guarantees and practical insights into its convergence behavior. The result demonstrates that the policy evaluation error can be systematically decomposed into two distinct components: the approximation error of the parametric function class and the algorithmic error arising from the stochastic approximation procedure. The first term, $\sup_{j\in\mathbb{I}_{m}}|v(\cdot,j)-v^{\xi^{*}}(\cdot,j)|_{C^{0}({\overline{{\cal D}}})}$, represents the inherent approximation capability of our chosen parametric family. This bias term is independent of the learning algorithm and reflects how well the optimal parameter $\xi^{*}$ can approximate the true value function within the selected function class. The second term, $Ci^{-\frac{\nu\rho_{\xi}}{2}}/\epsilon^{\rho_{\xi}/2}$, exhibits a polynomial decay with respect to the iteration number $i$ and vanishes asymptotically as the number of iterations increases, demonstrating the algorithm’s convergence to the optimal parameter configuration within the chosen function class.

### v2

@@ 2 Classical Optimal Switching Problem
$dX^{t,x,i,\alpha}_{s}=\mu(s,X^{t,x,i,\alpha}_{s},I_{s}^{t,i})ds+\sigma(s,X^{t,x,i,\alpha}_{s},I_{s}^{t,i})dW_{s},\quad s\in(t,T],$ (2.2)
@@
    The drift $\mu$ and volatility $\sigma$ have linear growth in $x$, that is, there exists a constant $C>0$ such that
    $|\mu(s,x,i)|+|\sigma(s,x,i)|\leq C(1+|x|)$
    for all $(s,x,i)\in[0,T]\times\mathds{R}^{n}\times\mathbb{I}_{m}$.
  • (iii)
@@ 3 Exploratory Formulation under Entropy Regularization
Given $M>0$, consider a smooth and non-decreasing cut-off function $\phi_{M}$ such that $\phi_{M}(x)=e^{x}$ for $x\leq M$, $\phi_{M}(x)\leq e^{M+1}$ for $x\in(M,M+1)$ and $\phi_{M}(x)=e^{M+1}$for $x\geq M+1$. Hence, $\phi_{M}$ is bounded and Lipschitz continuous. Denote ${\cal D}_{N}:=\{(t,x):(t,x)\in{\cal D},|x|<N\}$. First, we will solve the following PDE system: for $i\in\mathbb{I}_{m}$,
$\begin{cases}\displaystyle\frac{\partial V_{i}^{M,N}(t,x)}{\partial t}+\mathcal{L}^{i}_{x}V_{i}^{M,N}(t,x)+f(t,x,i)+\lambda\sum_{j\neq i}\phi_{M}\left(\frac{V^{M,N}_{j}(t,x)-V^{M,N}_{i}(t,x)-g_{ij}}{\lambda}\right)=0,\\[10.00002pt] \displaystyle\hfill(t,x)\in{\cal D}_{N},\\[5.0pt] \displaystyle V_{i}^{M,N}(t,x)=K(T-t)+h(x),\quad(t,x)\in\partial{\cal D}_{N},\end{cases}$ (3.7)
@@
Consequently, we can extract from $\{V_{i}^{N}(t,x)\}$ a subsequence converging uniformly in $\overline{{\cal D}}$ together with the first $x$ , $t$-derivatives and second $x$-derivatives to a limit function $V_{i}^{\lambda}$, which is a solution to the HJB system (3.4). Thus, we complete the proof of the theorem. ∎
@@
$V_{i}^{\lambda}(t,x)\geq\mathbb{E}_{t,x,i}\bigg[\int^{s}_{t}f(\ell,X_{\ell},I_{\ell})ds-\sum_{k=1}^{\infty}g_{\kappa_{k-1}\kappa_{k}}{\bf 1}_{\{\tau_{k}\leq s\}}+\lambda\int^{s}_{t}R({\bm{\pi}}_{\ell},I_{\ell})ds+V_{I_{s}}^{\lambda}(s,X_{s})\bigg].$ (3.17)
@@ 4 Policy Iteration and Convergence
$\begin{cases}\displaystyle\frac{\partial V_{i}^{n}(t,x)}{\partial t}+\mathcal{L}^{i}_{x}V_{i}^{n}(t,x)+f(t,x,i)+H_{i}({\bm{\pi}}^{n}_{i}(t,x),V_{1}^{n}(t,x),\cdots,V_{m}^{n}(t,x))=0,\\ \displaystyle V_{i}^{n}(T,x)=h(x).\end{cases}$ (4.1)
@@
By Proposition 3.3, the tuple $(V^{\lambda}_{1},\cdots,V_{m}^{\lambda})$ constitutes the optimal value function for the control problem (3.2). Since $(V^{n}_{1},\cdots,V^{n}_{m})$ is the value function corresponding to the strategy ${\bm{\pi}}^{n}(t,x)$ for the same problem, it follows that $V^{n}_{i}\leq V_{i}^{\lambda}$ for all $i\in\mathbb{I}_{m}$ and $n\geq 1$.
Given the uniform bound $V^{n}_{i}\leq V_{i}^{\lambda}$, the estimate (3.13), $V^{0}_{i}\in C^{0}(\overline{{\cal D}})$, and the policy iteration definition (4.3), we deduce that for any $n\geq 1$ and $i,j\in\mathbb{I}_{m}$, the function $\pi_{ij}^{n}(t,x)$ is bounded on $\overline{{\cal D}}$. For $n\geq 1$, proceeding with a truncation argument analogous to the proof of Lemma 3.1, we select a suitable constant $\tilde{K}$ and consider the following pair of PDE systems defined on the truncated domain ${\cal D}_{N}$:
$\begin{cases}\displaystyle\frac{\partial V_{i}^{n,N}(t,x)}{\partial t}+\mathcal{L}^{i}_{x}V_{i}^{n,N}(t,x)+f(t,x,i)\\ \displaystyle\quad+\,H_{i}({\bm{\pi}}^{n}_{i}(t,x),V_{1}^{n,N}(t,x),\cdots,V_{m}^{n,N}(t,x))=0,&(t,x)\in{\cal D}_{N},\\[10.00002pt] \displaystyle V_{i}^{n,N}(t,x)=\tilde{K}(T-t)+h(x),&(t,x)\in\partial{\cal D}_{N},\end{cases}$ (4.4)
and
$\begin{cases}\displaystyle\frac{\partial V_{i}^{n+1,N}(t,x)}{\partial t}+\mathcal{L}^{i}_{x}V_{i}^{n+1,N}(t,x)+f(t,x,i)\\ \displaystyle\quad+\,H_{i}({\bm{\pi}}^{n+1,N}_{i}(t,x),V_{1}^{n+1,N}(t,x),\cdots,V_{m}^{n+1,N}(t,x))=0,&(t,x)\in{\cal D}_{N},\\[10.00002pt] \displaystyle V_{i}^{n+1,N}(t,x)=\tilde{K}(T-t)+h(x),&(t,x)\in\partial{\cal D}_{N},\end{cases}$ (4.5)
where policy ${\bm{\pi}}^{n+1,N}$ is given by
$\pi_{ij}^{n+1,N}(t,x)=\exp\left(\frac{V^{n,N}_{j}(t,x)-g_{ij}-V^{n,N}_{i}(t,x)}{\lambda}\right),\penalty 10000\ i,j\in\mathbb{I}_{m},\penalty 10000\ j\neq i.$ (4.6)
As in Lemma 3.1, the bounded solutions satisfy $V^{n,N}_{i}\to V^{n}_{i}$ and $V^{n+1,N}_{i}\to V^{n+1}_{i}$ as $N\to\infty$ for each $i\in\mathbb{I}_{m}$.
Define the difference $\Delta_{i}^{n,N}(t,x):=V_{i}^{n+1,N}(t,x)-V_{i}^{n,N}(t,x)$ for $i\in\mathbb{I}_{m}$ and $(t,x)\in\overline{{\cal D}}_{N}$. Subtracting (4.4) from (4.5) yields the equation satisfied by $\Delta_{i}^{n,N}$:
$\frac{\partial\Delta_{i}^{n,N}(t,x)}{\partial t}+\mathcal{L}^{i}_{x}\Delta_{i}^{n,N}(t,x)$ $+H_{i}({\bm{\pi}}^{n+1,N}_{i}(t,x),V_{1}^{n+1,N}(t,x),\cdots,V_{m}^{n+1,N}(t,x))$
$-H_{i}({\bm{\pi}}^{n}_{i}(t,x),V_{1}^{n,N}(t,x),\cdots,V_{m}^{n,N}(t,x))=0,\quad(t,x)\in{\cal D}_{N},$ (4.7)
with boundary condition $\Delta_{i}^{n,N}(t,x)=0$ for $(t,x)\in\partial{\cal D}_{N}$. A key observation from (4.6) is that ${\bm{\pi}}_{i}^{n+1,N}(t,x)$ maximizes the Hamiltonian:
${\bm{\pi}}_{i}^{n+1,N}(t,x)=\underset{{\bm{\pi}}}{\mathrm{arg\,max\,}}H_{i}({\bm{\pi}},V_{1}^{n,N}(t,x),\cdots,V_{m}^{n,N}(t,x)).$ (4.8)
Using (4) and the maximizing property (4.8), we derive for $(t,x)\in{\cal D}_{N}$:
$\frac{\partial\Delta_{i}^{n,N}(t,x)}{\partial t}+\mathcal{L}^{i}_{x}\Delta_{i}^{n,N}(t,x)+\sum_{j\neq i}\pi_{ij}^{n+1,N}(t,x)\Delta_{j}^{n,N}(t,x)-\sum_{j\neq i}\pi_{ij}^{n+1,N}(t,x)\Delta_{i}^{n,N}(t,x)$
$=-H_{i}({\bm{\pi}}_{i}^{n+1,N}(t,x),V_{1}^{n+1,N}(t,x),\cdots,V_{m}^{n+1,N}(t,x))+\sum_{j\neq i}\pi_{ij}^{n+1,N}(t,x)(\Delta_{j}^{n,N}(t,x)-\Delta_{i}^{n,N}(t,x))$
$\quad+H_{i}({\bm{\pi}}^{n}_{i}(t,x),V_{1}^{n,N}(t,x),\cdots,V_{m}^{n,N}(t,x))$
$=H_{i}({\bm{\pi}}^{n}_{i}(t,x),V_{1}^{n,N}(t,x),\cdots,V_{m}^{n,N}(t,x))-H_{i}({\bm{\pi}}^{n+1,N}_{i}(t,x),V_{1}^{n,N}(t,x),\cdots,V_{m}^{n,N}(t,x))$
@@
Applying Theorem 1.3 in kusano1965first, we deduce that $\Delta^{n,N}_{i}(t,x)\geq 0$, that is, $V_{i}^{n+1,N}(t,x)\geq V_{i}^{n,N}(t,x)$, for all $i\in\mathbb{I}_{m}$ and $(t,x)\in\overline{{\cal D}}$. Passing to the limit as $N\to\infty$, we have $V_{i}^{n+1}(t,x)\geq V_{i}^{n}(t,x)$, which then completes the proof. ∎
@@
$\left|H_{i}({\bm{\pi}}_{i}^{n},V_{1}^{\lambda},\cdots,V_{m}^{\lambda})-H_{i}({\bm{\pi}}_{i}^{*},V_{1}^{\lambda},\cdots,V_{m}^{\lambda})\right|$
$=\Bigg|\sum_{j\neq i}\pi_{ij}^{n}(V_{j}^{\lambda}-g_{ij}-V_{i}^{\lambda})+\lambda\sum_{j\neq i}(\pi_{ij}^{n}-\pi_{ij}^{n}\log\pi_{ij}^{n})$
$\quad+\left(\sum_{j\neq i}\pi_{ij}^{*}(V_{j}^{\lambda}-g_{ij}-V_{i}^{\lambda})+\lambda\sum_{j\neq i}(\pi_{ij}^{*}-\pi_{ij}^{*}\log\pi_{ij}^{*})\right)\Bigg|$
@@
and $w_{i}^{n}(T,x)=0$ for $x\in\mathds{R}^{n}$. By virtue of Theorem 1.3 in kusano1965first, we deduce $w^{n}_{i}(t,x)\leq 0$. That is,
@@
$F^{n}(t)\leq\frac{(C^{*})^{n}T^{n}}{n!}\sup_{s\in[0,T]}F^{1}(s),\quad\forall t\in[0,T].$ (4.16)
Because $F^{1}(t)$ is bounded, let $C_{2}=C^{*}T$ and $C_{1}=\sup_{t\in[0,T]}F^{1}(t)$. Then we obtain that desired result. ∎
@@
It follows from (3.8), (3.13) and Assumption 2.2-(ii) that
$|V_{i}^{\lambda}(t,x)|\leq KT+K_{f,h}$ $=K_{f,h}T+\lambda\sup_{i\in\mathbb{I}_{m}}\left(\sum_{j\neq i}\exp\left(-\frac{g_{ij}}{\lambda}\right)\right)T+K_{f,h}$
$\leq K_{f,h}T+\lambda(m-1)T+K_{f,h}$
@@
On the other hand, it follows from the definition of upper and lower weak limits that $\overline{V}_{i}(t,x)\geq\underline{V}_{i}(t,x)$, for any $i\in\mathbb{I}_{m}$ and $(t,x)\in\overline{{\cal D}}$. Thus, $\overline{V}_{i}(t,x)=\underline{V}_{i}(t,x)$, and we denote by
$V^{*}_{i}(t,x):=\overline{V}_{i}(t,x)=\underline{V}_{i}(t,x)\quad\text{for}\penalty 10000\ i\in\mathbb{I}_{m},(t,x)\in\overline{{\cal D}}.$
@@ 5 Reinforcement Learning Algorithm
Recall that given a feedback strategy ${\bm{\pi}}(t,x)=(\pi_{ij}(t,x))_{i,j\in\mathbb{I}_{m}}$, the corresponding value function $(v_{1}^{{\bm{\pi}}},\cdots,v_{m}^{{\bm{\pi}}})$ satisfies the PDE system that for $i\in\mathbb{I}_{m}$,
@@
where the Hamiltomian $H_{i}$ is given by (4.2). For simplicity, we omit the superscript $\bm{\pi}$}} and denote the value function as:
@@
and denote by $I=(I_{t})_{t\geq 0}$ a continuous-time finite-state Markov chain with generator ${\bm{\pi}}=(\pi^{ij})_{i,j\in\mathbb{I}_{m}}$. Let us introduce the process $M=(M_{t})_{t\in[0,T]}$ given by
@@
Let ${\bm{\pi}}(t,x)=(\pi_{ij}(t,x))_{i,j\in\mathbb{I}_{m}}$ be a feedback strategy and $v(t,x,i)$ be the corresponding value function given by (5.2). Then the process $M=(M_{t})_{t\in[0,T]}$ given by (5.3) is a square-integrable martingale.
@@
$=M_{t^{\prime}}+\mathbb{E}\left[\int_{t^{\prime}}^{t}(D_{x}v(s,X_{s},I_{s}))^{\top}\sigma(s,X_{s},I_{s})dW_{s}\Big|\mathcal{F}_{t^{\prime}}\right]=M_{t^{\prime}}.$ (5.5)
@@
$\sup_{j\in\mathbb{I}_{m}}|v^{\xi_{i}}(\cdot,j)-v(\cdot,j)|_{C^{0}({\overline{{\cal D}}})}\leq\sup_{j\in\mathbb{I}_{m}}|v(\cdot,j)-v^{\xi^{*}}(\cdot,j)|_{C^{0}({\overline{{\cal D}}})}+\frac{C}{\epsilon^{\rho/2}}i^{-\frac{\nu\rho}{2}}.$ (5.14)
@@
$|\xi_{i}-\xi^{*}|\leq C\epsilon^{-\frac{1}{2}}i^{-\frac{\nu}{2}}$
@@
$\leq\sup_{j\in\mathbb{I}_{m}}|v^{\xi^{*}}(\cdot,)-v(\cdot,j)|_{C^{0}({\overline{{\cal D}}})}+\frac{C}{\epsilon^{\rho/2}}i^{-\frac{\nu\rho}{2}}.$
@@
Theorem 5.4 establishes a comprehensive error analysis for Algorithm 1, providing both theoretical guarantees and practical insights into its convergence behavior. The result demonstrates that the policy evaluation error can be systematically decomposed into two distinct components: the approximation error of the parametric function class and the algorithmic error arising from the stochastic approximation procedure. The first term, $\sup_{j\in\mathbb{I}_{m}}|v(\cdot,j)-v^{\xi^{*}}(\cdot,j)|_{C^{0}({\overline{{\cal D}}})}$, represents the inherent approximation capability of our chosen parametric family. This bias term is independent of the learning algorithm and reflects how well the optimal parameter $\xi^{*}$ can approximate the true value function within the selected function class. The second term, $Ci^{-\frac{\nu\rho}{2}}/\epsilon^{\rho/2}$, exhibits a polynomial decay with respect to the iteration number $i$ and vanishes asymptotically as the number of iterations increases, demonstrating the algorithm’s convergence to the optimal parameter configuration within the chosen function class.